# Understanding Stellar Mass-Metallicity and Size Relations in Simulated Ultra-Faint Dwarf Galaxies

### v1

@@ 3.1 Morphology and mass growth history
Fig. 4 presents the resulting stellar mass-halo mass (SMHM) relation for all progenitor halos at $z=6$ (circle) and the final UFD analogs at $z=0$ (cross). Progenitor halos that constitute the same UFD analog are represented by the same color. In addition, progenitors contributing less than 5% of the total stellar mass are also included. We find a positive correlation in the SMHM relation with a large scatter, particularly at $M_{\rm vir}\approx 10^{7}~{}\mbox{M}_{\odot}$, where stellar mass ranges in $M_{\star}\approx 10^{2}-10^{4}~{}\mbox{M}_{\odot}$. This spread could be attributed to the vulnerability of small-mass progenitor halos and their varying mass growth histories. For instance, if a progenitor halo of the same mass forms relatively late in a low-density region, it will have a shorter SFH, resulting in a lower stellar mass. This occurs because, after the onset of a late starburst, star formation will be immediately quenched by cosmic reionization, leaving no opportunity for subsequent star formation. Our SMHM relation for the UFD analogs at $z=0$ aligns well with other theoretical studies (e.g., Munshi et al., 2013; Wheeler et al., 2019; Applebaum et al., 2021), which demonstrate that stellar masses from $M_{\star}=10^{3}~{}\mbox{M}_{\odot}$ to a few $10^{4}~{}\mbox{M}_{\odot}$ are produced within a total halo mass range from $M_{\rm vir}=10^{8}~{}\mbox{M}_{\odot}$ to a few $10^{9}~{}\mbox{M}_{\odot}$.
@@
Figure 7: MDFs for our UFD analogs. To determine the average metallicity, $\rm\langle[Fe/H]\rangle$, and metallicity dispersion, $\sigma_{\rm[Fe/H]}$, for all stars (top panels), we utilize maximum likelihood estimation with a two-dimensional Gaussian likelihood function, similar to the method used for observational data (Fu et al., 2023). We also present the MDFs applying the observational limit at $\rm\langle[Fe/H]\rangle=-4$ in the bottom panels. Across all UFDs, the peaks fall within the range of $\rm[Fe/H]-3$ to $\rm[Fe/H]=-2$. Excluding stars with metallicities below $\rm[Fe/H]=-4$ results in an increase in average metallicity and a decrease in metallicity dispersion.
@@ 3.2.1 Stellar mass-metallicity relation
It is well-established that there is a positive correlation between the stellar mass and stellar metallicity of galaxies. This relationship is primarily due to more massive galaxies producing more stars and retaining more metals, which increases the average metallicity of their stars. Fig. 5 illustrates the resulting MZR for progenitor halos at $z\approx 6$ (circles) and the corresponding UFD analogs at $z=0$ (crosses). Our results show a positive correlation in the MZR, particularly among progenitor halos. Fig. 5 also demonstrates how stellar metallicity contributions from multiple progenitors can alter the final metallicities of the UFD analogs. For instance, although a progenitor with the highest $\rm\langle[Fe/H]\rangle$ value belonging to Halo5 (depicted in brown) shows $\rm\langle[Fe/H]\rangle=-2.0$ with $M_{\star}\sim 10^{4}~{}\mbox{M}_{\odot}$, the final $\rm\langle[Fe/H]\rangle$ value for Halo5 is reduced to $\rm\langle[Fe/H]\rangle=-2.71$ due to the contributions from other progenitor halos, lowering the overall $\rm\langle[Fe/H]\rangle$. A similar reduction is observed in Halo2, Halo4, and Halo5 as well. This implies that similar to the stellar mass assembly process, the metallicity of UFDs is not determined by the SFH of a single main progenitor. Instead, it is determined by the metallicity of stars formed in various environments of multiple progenitors.
To compare our results with other studies, Fig. 6 presents the derived MZR for galaxies based on both simulations and observations, covering a stellar mass range from $M_{\star}\approx 10^{2.5}~{}\mbox{M}_{\odot}$ to $10^{8.5}~{}\mbox{M}_{\odot}$. Observed galaxies are depicted in grey: circles denote UFDs and dwarf galaxies in the LG compiled by McConnachie (2012), while squares represent only MW UFDs reported by Fu et al. (2023). Simulated galaxies from cosmological simulations are shown in various colors and symbols. An increasing trend with rising stellar mass is evident in the classical dwarf regime ($M_{\star}\geq\rm 10^{5}~{}\mbox{M}_{\odot}$), showing alignment between observations and simulations. However, this correlation weakens for observed UFD galaxies that formed most of their stars before reionization, exhibiting a plateau rather than an increasing trend. Explaining this observed plateau has been challenging in theoretical work (e.g., Jeon et al., 2017; Wheeler et al., 2019; Agertz et al., 2020; Prgomet et al., 2022; Jeon et al., 2021b; Applebaum et al., 2021; Sanati et al., 2023), which shows that smaller-mass UFDs tend to have significantly lower average metallicity ($\rm\langle[Fe/H]\rangle\lesssim-3$).
@@
Some possible explanations have been suggested, including a metallicity-dependent IMF (e.g., Prgomet et al., 2022) or the consideration of Pop III star formation (e.g., Sanati et al., 2023). However, theoretical work still tends to show tension in the UFD regime, with overall $\rm\langle[Fe/H]\rangle$ values being 0.5-2.0 dex lower than observed values. This discrepancy is intriguing given the current ability to simulate UFD galaxies more realistically with exceptionally high resolutions ($m_{\rm gas}\approx 20-100~{}\mbox{M}_{\odot}$). Meanwhile, the observed plateau for UFDs might be due to tidal effects, as most of them are satellites of massive host galaxies (MW, M31), which could reduce their stellar mass without altering the average metallicity. However, utilizing cosmological simulations, Applebaum et al. (2021) found no significant difference in the average metallicity of simulated UFDs when comparing those with and without the host galaxy effect.
When comparing our results to other simulation studies, the values from Halo2 to Halo5 align closely with those having $\rm\langle[Fe/H]\rangle\approx-2.98$ to $-2.65$, about 1 dex higher than the values reported in other works (e.g., Wheeler et al., 2019; Applebaum et al., 2021; Sanati et al., 2023). Furthermore, our results show a smaller scatter in the $\rm\langle[Fe/H]\rangle$ values, with the lowest being $\rm\langle[Fe/H]\rangle\approx-2.98$, whereas other works report the lowest values as $\rm\langle[Fe/H]\rangle\lesssim-3.5$. We attribute these overall higher and narrower values, compared to other simulation studies, to the individual IMF sampling method adopted in this work.
The combined SN energy from the SSP stellar particle can exert excessive feedback on the surrounding medium, suppressing subsequent star formation for a period during which metals are too effectively dispersed, leading to the formation of subsequent stars with lower metallicity (e.g., Agertz et al., 2020; Applebaum et al., 2020). On the other hand, with individual IMF sampling, stars tend to form from gas that has been enriched by the previous SN explosion, immediately reflecting the gas metallicity shaped by that event. The impact of the IMF sampling method on the SFHs and resulting stellar metallicities of the simulated UFDs is also demonstrated in Jeon et al. (in preparation). Interestingly, Halo1 exhibits a higher average metallicity with $\rm\langle[Fe/H]\rangle=-2.19$. This is due to the unique characteristic of Halo1, which has only a single progenitor halo with a high $\rm\langle[Fe/H]\rangle$ value.
@@ 3.2.2 Metallicity distribution function
To investigate the detailed distribution of stellar metallicities, Fig. 7 presents MDFs of our simulated UFDs. The top panels display the MDF for all stars, while the bottom panels show the MDF for stars with $\rm[Fe/H]\geq-4$, reflecting the observational limit [^2] In Fu et al. (2023), they measured metallicity using the MESA Stellar Isochrones and Tracks (MIST) models (Choi et al., 2016; Dotter, 2016), where the metallicity grid is set from $\rm[Z/H]=-4$ to $\rm[Z/H]=-1$. (e.g., Fu et al., 2023). To determine the average metallicity ($\rm\langle[Fe/H]\rangle$) and metallicity dispersion ($\sigma_{\rm[Fe/H]}$) from the MDF, we use the same method applied to the observed UFD data in Fu et al. (2023), employing a maximum likelihood estimation with a two-dimensional Gaussian likelihood function (e.g., Walker et al., 2006). As shown in the top panels of Fig. 7, all the simulated galaxies exhibit a Gaussian-like distribution, with peaks in the range of $\rm-3\lesssim[Fe/H]\lesssim-2$ and dispersion values between $0.25\lesssim\sigma_{\rm[Fe/H]}\lesssim 0.76$. In the case of Halo1, the MDF has a narrower distribution than other halos, with the lowest dispersion of $\sigma_{\rm[Fe/H]}\rm=0.25$. This narrowness arises from having only one progenitor and star formation occurring in a single starburst event. On the other hand, since stars in the other simulated UFD analogs form through bursty star formation in multiple progenitor halos, the metallicity dispersion tends to increase. For example, Halo5, where stars formed in a total of 14 different progenitor halos prior to $z=6$, exhibits an increased metallicity dispersion, with $\sigma_{\rm[Fe/H]}\rm=0.75$.
One notable point in our results is the existence of low-metallicity stars with $\rm[Fe/H]<-4$. We observe that such low-metallicity stars primarily originate from external enrichment, where Pop II stars form first, without prior Pop III star formation, from gas contaminated by metals from nearby halos. The limited amount of metals from this external origin gives rise to low-metallicity stars. Meanwhile, stars formed via internal enrichment can achieve metallicities of $\rm[Fe/H]\geq-4$ through just one or two SN explosions. For example, in the prog1 of Halo4, a SN from a Pop III star with $m_{\star,\ \rm PopIII}\approx 60~{}\mbox{M}_{\odot}$ leads to the formation of a Pop II star with $\rm[Fe/H]=-3.5$. We find that only a few stars form through external metal enrichment, as most progenitor halos underwent internal enrichment for their initial ISM enrichment. Moreover, progenitor halos that experience external enrichment tend to have relatively small virial masses and are often located near more massive progenitor halos (e.g., Smith et al., 2015; Hicks et al., 2021).
@@
Figure 10: The procedure describing high-metallicity star formation during a starburst in prog2 of Halo5. Top Panel: The evolution of gas (solid lines) and metals (dotted lines) over cosmic time within spheres of radius 0.1 (cyan), 0.3 (grey), and 1 $R_{\rm vir}$ (black). Bottom Panel: Stellar metallicity as a function of star formation time is shown, focusing only on in situ stars. Blue crosses represent Pop III stars, Pop II SN stars by red crosses, and Pop II SSP stars that survive to $z=0$ are indicated by grey circles. After the formation of the first Pop II SSP star at $z=10.4$, a period of bursty star formation occurs, leading to the creation of high-metallicity stars from gas enriched by previous continuous star formation. However, subsequent cumulative SNe feedback significantly reduces the metal and gas mass, temporarily evacuating the gas within 0.1 $R_{\rm vir}$. This suppression halts further metal-enriched star formation in the progenitor.
@@
To figure out the general trend of MDFs for UFDs, Fig. 9 compares the composite MDFs for stars in all our simulated galaxies to those from observational data. The grey histogram displays stars observed in 13 MW UFDs as reported by Fu et al. (2023), while the black histogram includes stars in observed 26 UFDs in Local Group (LG), as detailed by Simon (2019). One evident difference in the composite MDFs is that, although both observed composite MDFs show a similar fraction of high-metallicity stars with $\rm[Fe/H]\geq-2.0$, reflecting the typical ratio of these stars in UFDs ($\sim$19%), our simulated galaxies lack such high-metallicity stars. This absence of high-metallicity stars may explain the discrepancy in the MZR between observations and simulations. Consequently, in the next section, we discuss the challenges in producing such high-metallicity stars in simulations of UFD analogs.
@@ 3.2.3 High-metallicity star formation
Here, we scrutinize the evolution of stellar metallicity by tracking the star-forming environment, gas, and metal reservoirs within a progenitor halo. Since the metallicity evolution in most progenitor halos exhibits a similar monotonically increasing pattern, we focus on prog2 of Halo5 as a representative example. In the top panel of Fig. 10, we display the mass evolution of gas (solid line) and metal (dotted line) within the progenitor halo as a function of cosmic time. We calculate the mass within 0.1 (cyan), 0.3 (grey), and 1 $R_{\rm vir}$ (black), respectively. This approach is taken because most stars form in the central region ($r_{\rm star}<0.3R_{\rm vir}$), where star formation is suppressed if there is insufficient gas. In the bottom panel, we show the metallicity of stars as a function of their formation time, focusing on the evolution of in situ stars. Blue crosses represent pop III stars, Pop II individual stars that explode as SNe are indicated by red crosses, and Pop II SSP stars that survive to $z=0$ are depicted in grey circles.
@@
We find the initial production of metals at $z\sim 10.7$, triggered by a Pop III SN. About 17 Myr later, a Pop II star forms at $z\sim 10.4$ with a metallicity of $\rm[Fe/H]=-4.7$, marking the onset of the first phase of continuous starburst. During this phase, the stellar metallicity increases monotonically from $\rm[Fe/H]=-3.1$ to $\rm[Fe/H]=-1.8$. This gradual increase is possible because highly dense regions are shielded from SNe, allowing the gas in these regions to be continuously enriched with metals. As shown in the top panel of Fig. 10, the metal mass increases during star formation, but its detailed evolution varies depending on the radius. The metal mass within the innermost sphere ($r<0.1R_{\rm vir}$) decreases rapidly due to SN feedback in the central region. This feedback also reduces the gas mass within all radii, with the gas in the innermost region ($r<0.1R_{\rm vir}$) even being evacuated completely. This effective suppression temporarily halts star formation in the halo. Star formation remains suppressed until a high-density region is reestablished at $z\sim 8.6$, about 100 Myrs after the first starburst phase.
During the second phase of continuous star formation at $z<9$, the initially formed in situ star in this phase has a metallicity of $\rm[Fe/H]\sim-3$, indicating that it forms from gas enriched by SN feedback of previous starburst, where the gas had diffused following the last star formation episode. The absence of SNe during this phase of continuous star formation leads to a monotonic increase in metallicity until a final Pop II star ends its life as SN, ending this star formation phase and resulting in temporary quenching. The last star in this halo, formed at $z\sim 6.7$, exerts strong feedback on the ISM combined with the effects of reionization, completely quenching any further star formation. Therefore, unlike massive galaxies, where stellar metallicity is generally expected to increase monotonically during prolonged star formation, high-metallicity stars in the simulated UFDs are likely to form during each starburst phase.

### v2

@@ 3.1 Morphology and mass growth history
Fig. 4 presents the resulting stellar mass-halo mass (SMHM) relation for all progenitor halos at $z=6$ (circle) and the final UFD analogs at $z=0$ (cross). Progenitor halos that constitute the same UFD analog are represented by the same color. In addition, progenitors contributing less than 5% of the total stellar mass are also included. We find a positive correlation in the SMHM relation with a large scatter, particularly at $M_{\rm vir}\approx 10^{7}~{}\mbox{M}_{\odot}$, where stellar mass ranges in $M_{\star}\approx 10^{2}-10^{4}~{}\mbox{M}_{\odot}$.
This spread could be attributed to the vulnerability of small-mass progenitor halos and their varying mass growth histories, which means their stellar mass is sensitively affected by nonlinear and stochastic processes during evolution. For instance, if a progenitor halo of the same mass forms relatively late in a low-density region, it will have a shorter SFH, resulting in a lower stellar mass. This occurs because, after the onset of a late starburst, star formation will be immediately quenched by cosmic reionization, leaving no opportunity for subsequent star formation. Our SMHM relation for the UFD analogs at $z=0$ aligns well with other theoretical studies (e.g., Munshi et al., 2013; Wheeler et al., 2019; Applebaum et al., 2021), which demonstrate that stellar masses from $M_{\star}=10^{3}~{}\mbox{M}_{\odot}$ to a few $10^{4}~{}\mbox{M}_{\odot}$ are produced within a total halo mass range from $M_{\rm vir}=10^{8}~{}\mbox{M}_{\odot}$ to a few $10^{9}~{}\mbox{M}_{\odot}$.
On the other hand, Revaz & Jablonka (2018) predicted $M_{\star}\sim 10^{5}~{}\mbox{M}_{\odot}$ in the halo mass within $[2\times 10^{8}-4\times 10^{9}]~{}\mbox{M}_{\odot}$. While showing a similar halo mass range to our UFD analogs, their stellar masses are roughly an order of magnitude higher. Given the similarities in feedback implementations—such as HI self-shielding and a UV background—the difference may partly stem from the delay time treatment. In our case, instantaneous SN feedback could lead to the early suppression of the star-forming region, compared to Revaz & Jablonka (2018), where the delayed feedback enables continuous star formation.
@@
[…]
@@ 3.2 Metallicity-related quantities
It is well-established that there is a positive correlation between the stellar mass and stellar metallicity of galaxies. This relationship is primarily due to more massive galaxies producing more stars and retaining more metals, which increases the average metallicity of their stars. Fig. 5 illustrates the resulting MZR for progenitor halos at $z\approx 6$ (circles) and the corresponding UFD analogs at $z=0$ (crosses). Our results show a positive correlation in the MZR, particularly among progenitor halos. Fig. 5 also demonstrates how stellar metallicity contributions from multiple progenitor halos can alter the final metallicities of the UFD analogs. For instance, although a progenitor halo with the highest $\rm\langle[Fe/H]\rangle$ value belonging to Halo5 (depicted in brown) shows $\rm\langle[Fe/H]\rangle=-2.0$ with $M_{\star}\sim 10^{4}~{}\mbox{M}_{\odot}$, the final $\rm\langle[Fe/H]\rangle$ value for Halo5 is reduced to $\rm\langle[Fe/H]\rangle=-2.71$ due to the contributions from other progenitor halos, lowering the overall $\rm\langle[Fe/H]\rangle$. A similar averaging effect is also observed in Halo2, Halo4, and Halo5. This implies that, similar to the stellar mass assembly process, the metallicity of UFDs is not determined by the single (main) progenitor. Instead, it is determined by multiple progenitor halos, within which stars exhibit a wide range of metallicities in different environments.
To compare our results with other studies, Fig. 6 presents the derived MZR for galaxies based on both simulations and observations, covering a stellar mass range from $M_{\star}\approx 10^{2.5}~{}\mbox{M}_{\odot}$ to $10^{8.5}~{}\mbox{M}_{\odot}$. Observed galaxies are depicted in grey: circles denote UFDs and dwarf galaxies in the LG compiled by McConnachie (2012), from which we use only those with metallicity measurements from spectroscopy, while squares represent MW UFDs reported by Fu et al. (2023). Simulated galaxies from cosmological simulations are shown in various colors and symbols. Throughout Fig. 5 and Fig. 6, we derive $\rm\langle[Fe/H]\rangle$ values using Pop II SSP star particles by the fitting method introduced in Sec. 3.2.2.
When comparing the results, an increasing trend with stellar mass is evident in the classical dwarf regime ($M_{\star}\geq\rm 10^{5}~{}\mbox{M}_{\odot}$), showing alignment between observations and simulations. However, this correlation weakens for observed UFD galaxies, exhibiting a plateau rather than an increasing trend. Explaining this observed plateau has been challenging in theoretical work (e.g., Jeon et al., 2017; Wheeler et al., 2019; Agertz et al., 2020; Prgomet et al., 2022; Jeon et al., 2021b; Applebaum et al., 2021; Sanati et al., 2023), which shows that smaller-stellar mass UFDs tend to have significantly lower average metallicity ($\rm\langle[Fe/H]\rangle\lesssim-3$).
@@
Some possible explanations have been suggested, including a metallicity-dependent IMF (e.g., Prgomet et al., 2022) or the consideration of Pop III star formation (e.g., Sanati et al., 2023). However, theoretical work still tends to show tension in the UFD regime, with overall $\rm\langle[Fe/H]\rangle$ values being 0.5-2.0 dex lower than observed values. This discrepancy is intriguing given the current ability to simulate UFD galaxies more realistically with high resolutions ($m_{\rm gas}\approx 20-100~{}\mbox{M}_{\odot}$). Meanwhile, the observed plateau for UFDs might be due to tidal effects, as most of them are satellites of massive host galaxies (MW, M31), which could reduce their stellar mass without altering the average metallicity. However, utilizing cosmological simulations, Applebaum et al. (2021) found no significant differences in the metallicity of simulated UFDs when comparing those with and without the host galaxy effect.
In comparison with other simulation studies, the $\rm\langle[Fe/H]\rangle$ values of Halo2 to Halo5 range from $\rm\langle[Fe/H]\rangle\approx-2.98$ to $-2.65$, which is about 1-1.5 dex higher than those reported in other works (e.g., Wheeler et al., 2019; Applebaum et al., 2021; Sanati et al., 2023). Furthermore, our results exhibit a smaller scatter in $\rm\langle[Fe/H]\rangle$ values, with the lowest being $\rm\langle[Fe/H]\rangle\approx-2.98$, whereas other studies yield the lowest values as $\rm\langle[Fe/H]\rangle\lesssim-3.5$. We attribute these overall higher and narrower values, compared to other simulation studies, to the individual IMF sampling method adopted in this work.
This can be understood as follows: if we use SSP scheme, the combined SNe energy from the SSP stellar particle can exert excessive feedback on the surrounding medium, suppressing subsequent star formation for a period during which metals are too effectively dispersed, leading to the formation of subsequent stars with lower metallicity (e.g., Agertz et al., 2020; Applebaum et al., 2020). On the other hand, with individual IMF sampling, stars tend to from gas that has been enriched by the previous SN explosion, immediately reflecting the gas metallicity shaped by that event. The impact of the IMF sampling method on the SFHs and resulting stellar metallicities of the simulated UFDs is also demonstrated in Jeon & Ko (2024).
Interestingly, Halo1 exhibits a higher average metallicity with $\rm\langle[Fe/H]\rangle=-2.19$. This uniqueness arises from two factors: first, Halo1 has a single progenitor, and second, this progenitor halo has high metallicity. At the progenitor halo scale, rather than the galaxy scale, this high metallicity is not particularly unusual. This is demonstrated by Fig. 5, where other halos also have progenitor halos with similarly high metallicities. However, for other halos, the average metallicity tends to be reduced as a result of mergers with low-metallicity progenitor halos during growth. In contrast, the average metallicity of Halo1 remains high because of its single high-metallicity progenitor with no further mergers. The properties of the progenitor halo that contribute to the high metallicity are discussed in Sec. 3.2.4.
@@ 3.2.2 Metallicity distribution function
To investigate the detailed distribution of stellar metallicities, Fig. 7 presents MDFs of our simulated UFDs. The top panels display the MDF for all stars, while the bottom panels show the MDF for stars with $\rm[Fe/H]\geq-4$, reflecting the observational limit [^3] In Fu et al. (2023), they measured metallicity using the MESA Stellar Isochrones and Tracks (MIST) models (Choi et al., 2016; Dotter, 2016), where the metallicity grid is set from $\rm[Z/H]=-4$ to $\rm[Z/H]=-1$. (e.g., Fu et al., 2023). To determine the average metallicity, $\langle\mathrm{[Fe/H]}\rangle$, and metallicity dispersion, $\sigma_{[\mathrm{Fe}/\mathrm{H}]}$, for Pop II SSP stars with masses below $8~{}~{}\mbox{M}_{\odot}$ (top panels), we utilize maximum likelihood estimation with a two-dimensional Gaussian likelihood function (e.g., Walker et al., 2006), which is the same method used for observational data in Fu et al. (2023).
As shown in the top panels of Fig. 7, all the simulated galaxies exhibit a Gaussian-like distribution, with peaks in the range of $\rm-3\lesssim[Fe/H]\lesssim-2$ and dispersion values between $0.25\lesssim\sigma_{\rm[Fe/H]}\lesssim 0.76$. In the case of Halo1, the MDF has a narrower distribution than other halos, with the lowest dispersion of $\sigma_{\rm[Fe/H]}\rm=0.25$. This narrowness arises from having only one progenitor and star formation occurring in a single starburst event. On the other hand, since stars in the other simulated UFD analogs form through bursty star formation in multiple progenitor halos, the metallicity dispersion tends to increase. For example, Halo5, where stars formed in a total of 14 different progenitor halos prior to $z=6$, exhibits an increased metallicity dispersion, with $\sigma_{\rm[Fe/H]}\rm=0.75$.
One notable point in our results is the existence of low-metallicity stars with $\rm[Fe/H]<-4$. We observe that such low-metallicity stars primarily originate from external enrichment, where Pop II stars form first, without prior Pop III star formation, from gas contaminated by metals from nearby halos. The limited amount of metals from the external halo gives rise to low-metallicity stars. In addition, these progenitor halos that experience external enrichment tend to have relatively small virial masses and are often located near more massive progenitor halos, as observed in other works (e.g., Smith et al., 2015; Hicks et al., 2021). Through all progenitor halos in our sets, few stars form through external metal enrichment, as most progenitor halos underwent internal enrichment for their initial ISM enrichment. For this case, stars with metallicities of $\rm[Fe/H]\geq-4$ form through just one or two SN explosions. For example, in the prog1 of Halo4, a SN from a Pop III star with $m_{\star,\ \rm PopIII}\approx 60~{}\mbox{M}_{\odot}$ leads to the formation of a Pop II star with $\rm[Fe/H]=-3.5$.
@@
Figure 10: The procedure describing high-metallicity star formation during a starburst in prog2 of Halo5. Top Panel: The evolution of gas (solid lines) and metals (dotted lines) over cosmic time within spheres of radius 0.1 (cyan), 0.3 (grey), and 1 $R_{\rm vir}$ (black). Middle Panel: Stellar metallicity as a function of star formation time is shown, focusing only on in situ stars. Blue crosses represent Pop III stars, Pop II SN stars by red crosses, and Pop II SSP stars that survive to $z=0$ are indicated by grey circles. After the formation of the first Pop II SSP star at $z=10.4$, a period of bursty star formation occurs, leading to the creation of high-metallicity stars from gas enriched by previous continuous star formation. However, subsequent cumulative SNe feedback significantly reduces the metal and gas mass, temporarily evacuating the gas within 0.1 $R_{\rm vir}$. This suppression halts further metal-enriched star formation in the progenitor. Bottom Panel: Star formation rates of each stellar type-Pop III (blue), Pop II individual star (red), and Pop II SSP, which is calculated binning 10 Myr.
@@
To figure out the general trend of MDFs for UFDs, Fig. 9 compares the composite MDFs for stars in all our simulated galaxies to those from observational data. The grey histogram displays stars observed in 13 MW UFDs as reported by Fu et al. (2023), while the black histogram includes stars in observed 26 UFDs in Local Group (LG), as detailed by Simon (2019). One evident difference in the composite MDFs is that, although both observed composite MDFs show a similar fraction of high-metallicity stars with $\rm[Fe/H]\geq-2.0$, reflecting the typical ratio of these stars in UFDs ($\sim$19%), our simulated galaxies lack such high-metallicity stars. This absence of high-metallicity stars may explain the discrepancy in the MZR between observations and simulations. In the next section, we discuss the challenges in producing such high-metallicity stars in simulations of UFD analogs.
@@ 3.2.3 High-metallicity star formation
Here, we scrutinize the evolution of stellar metallicity by tracking the star-forming environment, gas, and metal reservoirs within a progenitor halo. Since the metallicity evolution in most progenitor halos exhibits a similar monotonically increasing pattern, we focus on prog2 of Halo5 as a representative example. In the top panel of Fig. 10, we display the mass evolution of gas (solid line) and metal (dotted line) within the progenitor halo as a function of cosmic time. We calculate the mass within 0.1 (cyan), 0.3 (grey), and 1 $R_{\rm vir}$ (black), respectively. This approach is taken because most stars form in the central region ($r_{\rm star}<0.3\ R_{\rm vir}$), where star formation is suppressed if there is insufficient gas. In the middle panel, we show the metallicity of stars as a function of their formation time, focusing only on the evolution of in situ stars. Blue crosses represent Pop III stars; red crosses indicate Pop II individual stars that explode as SN; and grey circles denote Pop II SSP stars that survive to $z=0$. Finally, the bottom panel presents the star formation rate of these stars.
@@
We find the initial production of metals at $z\sim 10.7$, triggered by a Pop III SN. About 17 Myr later, a Pop II star forms at $z\sim 10.4$ with a metallicity of $\rm[Fe/H]=-4.7$, marking the onset of the first phase of starburst. During this phase, the stellar metallicity increases monotonically from $\rm[Fe/H]=-3.1$ to $\rm[Fe/H]=-1.8$. This increase in metallicity is possible by the localized and short timescale nature of SN feedback. Some dense regions, which have not yet formed stars, can be rapidly enriched by nearby SNe. However, due to their high density, they can withstand the injected energy without being disrupted, allowing star formation to proceed in their enriched state. As shown in the top panel of Fig. 10, metal mass within the innermost region ($r<0.1\ R_{\rm vir}$) initially increases due to metal ejection from Pop II SNe, and most of metals remain concentrated in this central region, although some diffuse outward. However, as SN energy continues to accumulate over time, it eventually disrupts the dense gas in the core, which causes a rapid outflow of metals from the center to the outer regions. This effective suppression temporarily halts star formation in the halo. Star formation remains suppressed until a high-density region is reestablished at $z\sim 8.6$, about 100 Myr after the first starburst phase.
In the second phase of starburst at $z<9$, shown in Fig. 10, the initially formed in situ star has a metallicity of $\rm[Fe/H]\sim-3$, indicating that it forms from gas enriched by SN feedback of previous starburst. During the period of 100 Myr between first and second starbursts, metals diffuse outward and disperse into extended regions, resulting in a decrease in the metallicity. This explains why the second starburst exhibits lower metallicities than the first starburst. Moreover, in the second starburst, individual stars that explode as SNe are sampled later in the phase, and this absence of early SNe leads to a monotonic decrease in metallicity during the starburst. The third starburst also features a late-forming SN progenitor at $z\sim 6.7$. This final SN feedback, combined with the effects of reionization, suppresses the ISM and completely quenches further star formation. Therefore, unlike in massive galaxies—where stellar metallicity typically increases monatonically over long timescale—high-metallicity stars in the UFD analogs are likely to form over short periods, specifically during each starburst phase.